OpenStudy (anonymous):

How do you factor x^2-3x+8?

OpenStudy (campbell_st):

you can't ... there are no factors of 8 that add to -3... both factors would be negative

OpenStudy (anonymous):

use completing the square

OpenStudy (anonymous):

you can still factor it...

OpenStudy (anonymous):

don't listen to cambell...it may not factor nicely but it is factorable

OpenStudy (campbell_st):

well the interesting thing with this question is there are no real roots ... ... just an observation

OpenStudy (anonymous):

still factorable, even if the roots are complex

OpenStudy (anonymous):

With the quadratics formula, it's possible, but I don't understand how to use it.

OpenStudy (anonymous):

just plug and chug.

OpenStudy (anonymous):

You can always add or subtract numbers to make it factor nicely. But if you do this then you will not be solve for =0 but instead some number,

OpenStudy (anonymous):

which is completing the square

OpenStudy (alexwee123):

wouldn't the roots be imaginary?

OpenStudy (campbell_st):

the formula is \[x=\frac{-b \pm \sqrt{b^2 -4ac}}{2a}\] in your question a = 1 b = -3 c = 8 substitute them and evaluate the problem will be inside the square root... you will get -23 this can be handled by the use of complex numbers...

OpenStudy (anonymous):

For example you can subtract 19 to both sides. x^2-3x+8=0 x^2-3x+8-19=-19 x^2-3x-10 = -19 (x-5)(x+2) = -19

OpenStudy (anonymous):

an easier number is half the coefficient of b, squared

OpenStudy (anonymous):

Can we keep the topic on math? Thanks. @Romero

OpenStudy (anonymous):

YOLO!!! ╭∩╮ - _ - ╭∩╮